# AI-based particle track identification in scintillating fibres read out with imaging sensors

### v1

@@ 2.1 Variational Autoencoders
Just like the standard autoencoder, a variational autoencoder (VAE) is composed of an encoder and a decoder which are trained to minimise a reconstruction error. However, as opposed to the standard autoencoder where the input is encoded as a single point, the variational autoencoder encodes the input as a distribution over the latent space, allowing the model to capture uncertainty in the data and creating a continuous and smooth latent space, among other benefits.
@@
The first term is equivalent to the reconstruction error and measures the reconstruction quality. The second term is the Kullback-Leibler Divergence. This ensures that the approximation $q_{\bm{\phi}}(\bm{z}|\bm{x})$ is close to the true posterior distribution $p_{\theta}(\bm{z}|\bm{x})$. [8, 9].
@@ 2.2 Anomaly Detection with VAE
However, the connection between anomalies and large reconstruction errors is not entirely clear [13]. Using VAEs has the advantage that the Kullback-Leibler Divergence term regularises the latent space by forcing it to follow a well-defined prior distribution. The model learns the distribution of latent variables instead of the variables itself. This means, we can label an event as anomalous if the probability of it being associated with the learned latent distribution is low. Therefore, instead of the reconstruction error, also the regularised latent space can be used for anomaly detection [4].
@@ 3 Anomaly Detection with VAE
This section details the VAE model architecture, training process, and its application for anomaly detection.
@@ 3.1 VAE Architecture and Training Details
The loss function $\mathcal{L}$ combines the Binary Cross-Entropy (BCE) loss and Kullback-Leibler Divergence (KLD).
@@
The VAE was implemented using PyTorch 1.14.0 [22] with Python 3.8.10 and trained on an NVIDIA V100 GPU with a batch size of 256 for 50 epochs. The model parameters were optimised using the AdamW optimiser with a learning rate of $1\times 10^{-3}$. To ensure a smooth training process, we employed a learning rate schedule that consisted of a linear warm-up phase followed by cosine annealing. The warm-up phase was applied during the first epoch, after which a cosine annealing schedule was used for the remaining epochs. This combined scheduling approach facilitated stable initial convergence while enabling efficient exploration of the loss landscape during later stages of training. The $\mathrm{BG}$ dataset consists of $5,012,736$ frames, split in the following way: $60\%$ for training, $10\%$ for validation, and $30\%$ for testing. The training and validation losses are shown in Fig. 2.
@@ 3.2 Anomaly Detection
To identify such signal frames in the ${}^{90}\mathrm{Sr+BG}$ dataset, we utilise the loss values from our VAE. Figure 4 illustrates the distribution of BCE and KLD loss values for the test set of the $\mathrm{BG}$ and the ${}^{90}\mathrm{Sr+BG}$ dataset. The plots indicate the divergence points, where the loss distributions of the ${}^{90}\mathrm{Sr+BG}$ data start to exceed those of the BG test set. Both these divergence points and the 98th percentile of the $\mathrm{BG}$ test set, which is also marked in the plots, are used to identify anomalies.
@@
Figure 4: Comparison of loss distributions for the Variational Autoencoder (VAE) model. The plot on the left shows the distribution of the Binary Cross-Entropy (BCE) reconstruction loss, while the right shows the distribution of the Kullback-Leibler Divergence (KLD) loss. In both plots, the blue curve represents the loss distribution for the $\mathrm{BG}$ test set, and the red curve represents the loss distribution for the ${}^{90}\mathrm{Sr+BG}$ data. The vertical dashed green lines indicate the divergence points, where the loss distributions of the ${}^{90}\mathrm{Sr+BG}$ data begin to exceed those of the $\mathrm{BG}$ test set. The vertical dashed-dotted purple lines mark the 98th percentile of the loss distribution for the $\mathrm{BG}$ test set, serving as a threshold to identify outliers or regions of interest where the loss values are notably high.
An example of our anomaly detection method is presented in Fig. 5, where we utilise the 98th percentile of the BCE and KLD losses from the $\mathrm{BG}$ test set to select frames in both the $\mathrm{BG}$ test set and the ${}^{90}\mathrm{Sr+BG}$ dataset. For the $\mathrm{BG}$ sample, the separation appears completely random to the naked eye (Fig. 5(a)). On the other hand, for the ${}^{90}\mathrm{Sr+BG}$, the figure illustrates the VAE’s ability to extract signal frames, where the pattern of short electron tracks traversing from top to bottom is distinctly visible, in contrast to the remaining candidates, where the summed frames resemble random noise (Fig. 5(b)).
@@
Figure 5: Summed frames within the fibre bundle region, separated based on whether they are considered anomalies (either BCE or KLD loss exceeded the 98th percentile on the $\mathrm{BG}$ test set) or not.
@@
Table 1 shows the correlation between the number of counts and selected frames. The percentage of selected frames as anomaly candidates is higher in the ${}^{90}\mathrm{Sr+BG}$ sample compared to the $\mathrm{BG}$ test set for every number of counts per frame. In most cases, it is significantly higher or at least the same. We can conclude that the selection of a frame is not due to its higher number of counts but rather due to the signal event it likely contains. Of course, depending on the goal of the analysis, one can lower the threshold used in the BCE and KLD losses to select anomalies to avoid missing signals, but at the cost of allowing more background frames (high selection efficiency) or minimise the background at the risk of missing some signals (high selection purity). Additional results using different configurations of loss terms and thresholds are provided in Appendix A. Both the KLD and BCE loss terms prove useful for anomaly detection.
@@
Frames Selected % Frames Selected %
4 23396 104 0.44 153066 2944 1.92
5 4830 367 7.60 44146 13630 30.87
6 828 424 51.21 14397 12457 86.52
7 109 95 87.16 7521 7494 99.64
8 14 14 100.00 5768 5768 100.00
9 1 1 100.00 5159 5159 100.00
@@ 3.3 Fast inference
Encoder CPU 3.69 $\pm$ 1.23 1.82 $\pm$ 0.58 1.28 $\pm$ 0.66 0.69 $\pm$ 0.38 0.35 $\pm$ 0.16 0.22 $\pm$ 0.11 0.10 $\pm$ 0.12
GPU 1.60 $\pm$ 7.50 0.72 $\pm$ 0.93 0.38 $\pm$ 0.13 0.21 $\pm$ 0.15 0.10 $\pm$ 0.05 0.05 $\pm$ 0.05 0.01 $\pm$ 0.02
Decoder CPU 19.05 $\pm$ 1.86 15.73 $\pm$ 0.72 15.42 $\pm$ 0.70 15.35 $\pm$ 0.46 15.08 $\pm$ 0.45 13.94 $\pm$ 4.34 6.94 $\pm$ 7.58
GPU 0.46 $\pm$ 6.96 0.31 $\pm$ 0.89 0.13 $\pm$ 0.11 0.09 $\pm$ 0.14 0.09 $\pm$ 0.05 0.08 $\pm$ 0.05 0.04 $\pm$ 0.05
KLD CPU 0.17 $\pm$ 0.02 0.09 $\pm$ 0.01 0.05 $\pm$ 0.01 0.03 $\pm$ 0.00 0.02 $\pm$ 0.00 0.01 $\pm$ 0.00 0.00 $\pm$ 0.00
GPU 0.25 $\pm$ 1.43 0.10 $\pm$ 0.01 0.05 $\pm$ 0.01 0.03 $\pm$ 0.00 0.01 $\pm$ 0.00 0.01 $\pm$ 0.00 0.00 $\pm$ 0.00
BCE CPU 0.90 $\pm$ 0.29 0.54 $\pm$ 0.16 0.30 $\pm$ 0.06 0.16 $\pm$ 0.01 0.09 $\pm$ 0.01 0.07 $\pm$ 0.02 0.03 $\pm$ 0.03
GPU 0.13 $\pm$ 0.49 0.06 $\pm$ 0.01 0.03 $\pm$ 0.00 0.02 $\pm$ 0.00 0.01 $\pm$ 0.00 0.00 $\pm$ 0.00 0.00 $\pm$ 0.00
Memory allocation GPU 0.16 $\pm$ 0.04 0.11 $\pm$ 0.02 0.09 $\pm$ 0.01 0.07 $\pm$ 0.01 0.05 $\pm$ 0.00 0.04 $\pm$ 0.01 0.02 $\pm$ 0.02
Total CPU 23.81 $\pm$ 1.51 18.18 $\pm$ 0.74 17.05 $\pm$ 0.58 16.24 $\pm$ 0.57 15.53 $\pm$ 0.48 14.23 $\pm$ 4.43 7.07 $\pm$ 7.72
GPU 2.60 $\pm$ 2.50 1.31 $\pm$ 0.08 0.69 $\pm$ 0.05 0.41 $\pm$ 0.03 0.26 $\pm$ 0.01 0.18 $\pm$ 0.06 0.08 $\pm$ 0.08
@@ 4 Conclusion
Scintillating fibres read out with imaging sensors offer an optimal configuration for high spatial resolution particle tracking. Given the growing success of machine learning techniques for anomaly detection, we proposed using a variational autoencoder (VAE) to differentiate between signal and background frames. In this study, we demonstrated that a VAE, trained solely on background frames, can serve as an effective and fast selection algorithm. We used the Binary Cross-Entropy (BCE) and the Kullback-Leibler Divergence (KLD) losses as anomaly scores to identify signal frames. The results showed that the frames selected indeed contained signal events. Furthermore, we tested the same method on background frames and found that significantly fewer frames were selected, confirming the effectiveness of the VAE with both BCE and KLD losses as anomaly detection metrics. Although we have not yet tested this approach with finer fibres, we hypothesise that the VAE would further improve its performance, given the lower number of dark counts and the more refined pattern left by a particle, which would be easier and faster for the method to detect. In future work, we plan to explore its integration as a real-time anomaly detection method into the hardware.

### v3

@@ 2.1 Variational Autoencoders
Just like the standard autoencoder, a VAE is composed of an encoder and a decoder which are trained to minimise a reconstruction error. However, as opposed to the standard autoencoder where the input is encoded as a single point, the VAE encodes the input as a distribution over the latent space, allowing the model to capture uncertainty in the data and creating a continuous and smooth latent space, among other benefits.
@@
The first term is equivalent to the reconstruction error and measures the reconstruction quality. The second term is the Kullback-Leibler Divergence (KLD). This ensures that the approximation $q_{\bm{\phi}}(\bm{z}|\bm{x})$ is close to the true posterior distribution $p_{\theta}(\bm{z}|\bm{x})$. [8, 9].
@@ 2.2 Anomaly Detection with VAE
However, the connection between anomalies and large reconstruction errors is not entirely clear [13]. Using VAEs has the advantage that the KLD term regularises the latent space by forcing it to follow a well-defined prior distribution. The model learns the distribution of latent variables instead of the variables itself. This means, we can label an event as anomalous if the probability of it being associated with the learned latent distribution is low. Therefore, instead of the reconstruction error, also the regularised latent space can be used for anomaly detection [4].
@@ 3 Anomaly Detection with VAE
This section details the datasets along with a standard method for identifying tracks. Besides, the VAE model architecture, training process, and its application for anomaly detection are also presented.
3.1 Experimental Setup and Data
Figure 2: (a) SciFi setup used for the acquisition of the data samples. Scintillating fibres are independently coupled to SiPM on their left side end, while on the opposite side they are bundled together and coupled to a single SwissSPAD2 sensor. (b) Close-up photograph of the face of $4\times 4$ scintillating fibre bundle. (c) Coupling of the scintillating fibre bundle with SwissSPAD2. Black glob top was used to protect the wire bonding around the active area. Taken from [3].
The dataset analysed in this study is the same as in Ref. [3]. Sixteen square scintillating fibres [20] with a diameter of 1 mm were arranged into a 4×4 bundle and coupled to SwissSPAD2 [21]. On top of the bundle of fibres, a 185 kBq-90Sr source could be placed; see Fig. 2. The electrons from the source then travel through the fibres, creating scintillation photons, which are subsequently detected by the SPAD. The setup was placed in a dark chamber to limit noise due to stray light.
A first sample, without the 90Sr source present, subsequently called the background ($\mathrm{BG}$) sample, consists of $5,012,736$ frames. A single frame contains all detected counts in a 1 $\mu s$ time window. Counts in these frames are due to sensor noise, known as dark counts. A second sample, this time with the 90Sr source placed centrally on top of the fibres, contains the same number of frames and is subsequent called the signal sample (${}^{90}\mathrm{Sr+BG}$). Counts in this sample can contain both dark counts and detected scintillation photons.
Figure 3: Distribution of number of counts per 1 $\mu s$ frame corresponding to the fibre bundle region on SwissSPAD2. The hollow red markers show the distribution obtained from data collected without 90Sr using SwissSPAD2, known as the background ($\mathrm{BG}$) data. The grey histogram is a Poisson distribution with $\lambda$ equal to mean counts of the $\mathrm{BG}$ data. The solid blue dots show the distribution of the number of counts per frame taken with the 90Sr source positioned above the fibre bundle. Adjusted from [3].
In [3], a simple cutoff in the number of detected counts was used to differentiate between signal- and background-like frames in the ${}^{90}\mathrm{Sr+BG}$ sample. The cutoff used was ten counts, the highest number of counts detected in the background sample. Figure 3 shows that in the intermediate region between five and ten counts, potential signal frames are present. Identifying signal-like frames in this region requires more complex analysis steps.
3.2 Standard Method for Identifying Electron Tracks
Due to the lack of available studies performing systematic track selection for scintillating fibre detectors read out by imaging sensors, we describe a standard method for identifying electron tracks inspired by the approach used in Ref. [3]. This reference method is designed to detect vertical track-like events using pixel-to-fibre mapping, requiring electrons to traverse a minimum number of fibres while accumulating a specified number of counts per track and per fibre (see Fig. 4). This selection strategy leverages the expected behaviour of electrons from a 90Sr source as they interact with the scintillating fibres, producing detectable photon signals. To define a valid track, the method enforces a minimum count threshold per fibre and requires a continuous trajectory across multiple fibres. Frames that fail to meet these criteria are discarded as background events. This method aims to distinguish true electron-induced signals from random noise and dark counts, thereby improving the reliability of track identification in experimental data.
Figure 4: The standard track selection process, following the approach described in Ref. [3], is illustrated for three example frames. The red lines indicate the fibre contours, while the filled squares highlight fibres with at least one photon count. The left frame is identified as a signal event, as it contains detected photons on at least three vertically down-going fibres. Similarly, the central frame is classified as signal due to the presence of a diagonal down-going track. In contrast, the right frame is categorised as noise, as it lacks a discernible down-going pattern.
It is important to note that this selection method was originally applied to the entire dataset, which includes frames with a wide range of detected counts, often exceeding 10. The effectiveness of the approach was primarily demonstrated in high-count regions, where track identification is more reliable due to the increased signal-to-noise ratio.
@@ 3.3 VAE Architecture and Training Details
The loss function $\mathcal{L}$ combines the Binary Cross-Entropy (BCE) loss and KLD.
@@
The VAE was implemented using PyTorch 1.14.0 [24] with Python 3.8.10 and trained on an NVIDIA V100 GPU with a batch size of 256 for 50 epochs. The model parameters were optimised using the AdamW optimiser with a learning rate of $1\times 10^{-3}$. To ensure a smooth training process, we employed a learning rate schedule that consisted of a linear warm-up phase followed by cosine annealing. The warm-up phase was applied during the first epoch, after which a cosine annealing schedule was used for the remaining epochs. This combined scheduling approach facilitated stable initial convergence while enabling efficient exploration of the loss landscape during later stages of training. The $\mathrm{BG}$ dataset was split in the following way: $60\%$ for training, $10\%$ for validation, and $30\%$ for testing. The training and validation losses are shown in Fig. 5.
@@ 3.4 Anomaly Detection
To identify such signal frames in the ${}^{90}\mathrm{Sr+BG}$ dataset, we utilise the loss values from our VAE. Figure 7 illustrates the distribution of BCE and KLD loss values for the test set of the $\mathrm{BG}$ and the ${}^{90}\mathrm{Sr+BG}$ dataset. The plots indicate the divergence points, where the loss distributions of the ${}^{90}\mathrm{Sr+BG}$ data start to exceed those of the BG test set. Both these divergence points and the 98th percentile of the $\mathrm{BG}$ test set, which is also marked in the plots, are used to identify anomalies. The 98th percentile was chosen empirically to ensure anomalies detected with this threshold are well beyond the divergence point while balancing signal retention and background rejection. Tests with alternative thresholds from the 95th percentile yielded consistent results, demonstrating the method’s robustness to small variations in the cutoff.
@@
Figure 7: Comparison of loss distributions for the Variational Autoencoder (VAE) model. The plot on the left shows the distribution of the Binary Cross-Entropy (BCE) reconstruction loss, while the right shows the distribution of the Kullback-Leibler Divergence (KLD) loss. In both plots, the blue curve represents the loss distribution for the $\mathrm{BG}$ test set, and the red curve represents the loss distribution for the ${}^{90}\mathrm{Sr+BG}$ data. The vertical dashed green lines indicate the divergence points, where the loss distributions of the ${}^{90}\mathrm{Sr+BG}$ data begin to exceed those of the $\mathrm{BG}$ test set. The vertical dashed-dotted purple lines mark the 98th percentile of the loss distribution for the $\mathrm{BG}$ test set, serving as a threshold to identify outliers or regions of interest where the loss values are notably high. The distributions are computed by evaluating the BCE and KLD losses for each individual frame in the dataset. The x-axis represents the per-frame loss values, and the y-axis represents the probability density function (PDF) estimated using Kernel Density Estimation (KDE). This visualisation allows us to identify points where the loss values of the ${}^{90}\mathrm{Sr+BG}$ dataset start deviating from the $\mathrm{BG}$ test set, motivating the selection of thresholds for anomaly detection.
An example of our anomaly detection performance is presented in Fig. 8, which compares the standard method described in Sec. 3.2 with the VAE. Frames are selected in both the $\mathrm{BG}$ test set and the ${}^{90}\mathrm{Sr+BG}$ dataset. The VAE utilises the 98th percentile of the BCE and KLD losses from the $\mathrm{BG}$ test set to select frames. For the $\mathrm{BG}$ sample, the separation appears completely random to the naked eye for both cases (Fig. 8(a)). On the other hand, for the ${}^{90}\mathrm{Sr+BG}$, the figure illustrates the VAE’s ability to extract signal frames, where the pattern of short electron tracks traversing from top to bottom is distinctly visible, in contrast to the remaining candidates, where the summed frames resemble random noise (Fig. 8(b)). The selection made by the standard method seems more arbitrary, and clear signal patters seem to remain in the non-selected frames. This behaviour is also evident in Fig. 9, where the total photon count after excluding the VAE-selected candidates exhibits a significantly flatter distribution compared to the result obtained by subtracting the standard method’s track candidates, meaning that it rejects dark count events more efficiently. This observation indicates that the VAE is more effective at capturing signal events.
@@
Figure 8: Summed frames within the fibre bundle region, separated based on whether they are selected or not with the Standard Method (Sec. 3.2) and with the VAE (either BCE or KLD loss exceeded the 98th percentile on the $\mathrm{BG}$ test set).
Figure 9: Photon count for each pixel row after summing all the frames for different scenarios on the ${}^{90}\mathrm{Sr+BG}$ sample.
@@
Table 1 shows the correlation between the number of counts and selected frames for both methods. For the VAE, the percentage of selected frames as anomaly candidates is higher in the ${}^{90}\mathrm{Sr+BG}$ sample compared to the $\mathrm{BG}$ test set for every number of counts per frame. In most cases, it is significantly higher or at least the same. We can conclude that the selection of a frame is not due to its higher number of counts but rather due to the signal event it likely contains. Of course, depending on the goal of the analysis, one can lower the threshold used in the BCE and KLD losses to select anomalies to avoid missing signals, but at the cost of allowing more background frames (high selection efficiency) or minimise the background at the risk of missing some signals (high selection purity). Additional results using different configurations of loss terms and thresholds are provided in Appendix A. Both the KLD and BCE loss terms prove useful for anomaly detection. In contrast, the standard method, which was designed for frames across all count ranges (including those with more than ten counts), exhibits similar selection rates for both the $\mathrm{BG}$ test set and ${}^{90}\mathrm{Sr+BG}$ sample in the 5-9 counts per frame range. This highlights the challenge of anomaly detection in low-count scenarios, where traditional selection criteria fail to distinguish between background and signal frames effectively. Unlike the VAE-based approach, which adapts to subtle distribution shifts in the anomaly detection process, the standard method struggles to capture meaningful differences, leading to a more uniform selection rate across both samples.
Although standard anomaly detection metrics such as precision, recall, or AUC are commonly used to evaluate classification performance, their application in this study is limited by the absence of ground truth labels. Unlike simulated datasets where the true classification of each frame is known, the experimental nature of our dataset does not allow a precise quantification of detection performance. However, the results presented in this section demonstrate that our method successfully isolates signal events, as summing anomalous and non-anomalous frames separately reveals a clear track pattern for the detected anomalies.
@@
Frames VAE Standard Frames VAE Standard
Selected % Selected % Selected % Selected %
4 23396 104 0.44 1865 7.97 153066 2944 1.92 12400 8.10
5 4830 367 7.60 813 16.83 44146 13630 30.87 6903 15.64
6 828 424 51.21 220 26.57 14397 12457 86.52 3020 20.98
7 109 95 87.16 46 42.20 7521 7494 99.64 1710 22.74
8 14 14 100.00 4 28.57 5768 5768 100.00 1469 25.47
9 1 1 100.00 1 100.00 5159 5159 100.00 1618 31.36
@@ 3.5 Fast inference
Encoder CPU 3.7 $\pm$ 1.2 1.8 $\pm$ 0.6 1.3 $\pm$ 0.7 0.7 $\pm$ 0.4 0.4 $\pm$ 0.2 0.2 $\pm$ 0.1 0.1 $\pm$ 0.1
GPU 1.6 $\pm$ 7.5 0.7 $\pm$ 0.9 0.4 $\pm$ 0.1 0.2 $\pm$ 0.2 0.1 $\pm$ 0.1 0.1 $\pm$ 0.1 0.0 $\pm$ 0.0
Decoder CPU 19.1 $\pm$ 1.9 15.7 $\pm$ 0.7 15.4 $\pm$ 0.7 15.4 $\pm$ 0.5 15.1 $\pm$ 0.5 14.0 $\pm$ 4.3 7.0 $\pm$ 7.6
GPU 0.5 $\pm$ 7.0 0.3 $\pm$ 0.9 0.1 $\pm$ 0.1 0.1 $\pm$ 0.1 0.1 $\pm$ 0.1 0.1 $\pm$ 0.1 0.0 $\pm$ 0.1
KLD CPU 0.2 $\pm$ 0.0 0.1 $\pm$ 0.0 0.1 $\pm$ 0.0 0.0 $\pm$ 0.0 0.0 $\pm$ 0.0 0.0 $\pm$ 0.0 0.0 $\pm$ 0.0
GPU 0.3 $\pm$ 1.4 0.1 $\pm$ 0.0 0.1 $\pm$ 0.0 0.0 $\pm$ 0.0 0.0 $\pm$ 0.0 0.0 $\pm$ 0.0 0.0 $\pm$ 0.0
BCE CPU 0.9 $\pm$ 0.3 0.5 $\pm$ 0.2 0.3 $\pm$ 0.1 0.2 $\pm$ 0.0 0.1 $\pm$ 0.0 0.1 $\pm$ 0.0 0.0 $\pm$ 0.0
GPU 0.1 $\pm$ 0.5 0.1 $\pm$ 0.0 0.0 $\pm$ 0.0 0.0 $\pm$ 0.0 0.0 $\pm$ 0.0 0.0 $\pm$ 0.0 0.0 $\pm$ 0.0
Memory allocation GPU 0.2 $\pm$ 0.0 0.1 $\pm$ 0.0 0.1 $\pm$ 0.0 0.1 $\pm$ 0.0 0.1 $\pm$ 0.0 0.0 $\pm$ 0.0 0.0 $\pm$ 0.0
Total CPU 23.8 $\pm$ 1.5 18.2 $\pm$ 0.7 17.1 $\pm$ 0.6 16.2 $\pm$ 0.6 15.5 $\pm$ 0.5 14.2 $\pm$ 4.4 7.1 $\pm$ 7.7
GPU 2.6 $\pm$ 2.5 1.3 $\pm$ 0.1 0.7 $\pm$ 0.1 0.4 $\pm$ 0.0 0.3 $\pm$ 0.0 0.2 $\pm$ 0.1 0.1 $\pm$ 0.1
@@ 4 Conclusion
Scintillating fibres read out with imaging sensors offer an optimal configuration for high spatial resolution particle tracking. Given the growing success of machine learning techniques for anomaly detection, we proposed using a VAE to differentiate between signal and background frames. In this study, we demonstrated that a VAE, trained solely on background frames, can serve as an effective and fast selection algorithm. We used the BCE and the KLD losses as anomaly scores to identify signal frames. The results showed that the frames selected indeed contained signal events. Furthermore, we tested the same method on background frames and found that significantly fewer frames were selected, confirming the effectiveness of the VAE with both BCE and KLD losses as anomaly detection metrics. In addition, a comparison with a standard selection method demonstrated that the VAE approach significantly outperforms traditional techniques, particularly in challenging low-count scenarios. Although we have not yet tested this approach with finer fibres, we hypothesise that the VAE would further improve its performance, given the lower number of dark counts and the more refined pattern left by a particle, which would be easier and faster for the method to detect. In future work, we plan to explore its integration as a real-time anomaly detection method into the hardware.